# Robust Survival Analysis with Adversarial Regularization

### v1

@@ Abstract
Survival Analysis (SA) is about modeling the time for an event of interest to occur, which has important applications in many fields, including medicine, defense, finance, and aerospace. Recent work has demonstrated the benefits of using Neural Networks (NNs) to capture complicated relationships in SA. However, the datasets used to train these models are often subject to uncertainty (e.g., noisy measurements, human error), which we show can substantially degrade the performance of existing techniques. To address this issue, this work leverages recent advances in NN verification to provide new algorithms for generating fully-parametric survival models that are robust to such uncertainties. In particular, we introduce a robust loss function for training the models and use CROWN-IBP regularization to address the computational challenges with solving the resulting Min-Max problem. To evaluate the proposed approach, we apply relevant perturbations to publicly available datasets in the SurvSet repository and compare survival models against several baselines. We empirically show that Survival Analysis with Adversarial Regularization (SAWAR) method on average ranks best for dataset perturbations of varying magnitudes on metrics such as Negative Log Likelihood (NegLL), Integrated Brier Score (IBS), and Concordance Index (CI), concluding that adversarial regularization enhances performance in SA.
@@
Survival Analysis, Adversarial Robustness, Neural Networks, Calibration, CROWN-IBP
publicationid: pubid: $0$–$00$ 2024 IEEE
@@ I Introduction
Survival Analysis (SA) is a highly active research field [1, 2] used to model the time before an event happens. For example, SA has been used extensively in medicine to quantify the effectiveness of various treatments as part of studies for cancer patients and drug trials [3, 4, 5]. While early approaches (dating back to the 1950s) predominantly used linear parametric models [6, 7], recent methods often leverage Neural Networks (NNs) to model complicated time-to-event data [8].
A key open challenge in this field is in developing models that are robust to uncertainty and variability in the datasets. Successful generalization of machine learning models over unseen datasets is crucial for clinical applications, as efficacy of developed drugs or treatments should be demonstrated over diverse patient demographics, institutions, and data acquisition settings [9, 10, 11]. Data uncertainty could be caused by human errors (e.g., notoriously poor handwriting of physicians [12], ineffective medical record keeping [13]), or inherent noise during transmissions or data collection [14]. However, many approaches assume perfect measurement of the covariates, ignoring aleatoric uncertainty [15]. In SA, the aleatoric uncertainty is only associated with the noise of the log of the measured time-to-event [16, 6]. Moreover, several studies have shown that adversaries can influence the final model to be over-/under-confident by purposefully poisoning studies or otherwise disrupting the data collection activities [17, 18]. Despite these known failure modes, and coupled with the sensitivity of NNs to input uncertainties, addressing the robustness of NNs against $\epsilon$-perturbations trained specifically for fully-parametric SA remains unexplored.
@@
  • identification of the sensitivity of existing SA methods, such as Deep Regularized Accelerated Failure Time (DRAFT) [19], to covariate perturbations via evaluating open-source datasets with a wide range of metrics related to predictive accuracy, calibration, and population curve estimation.
  • our approach, Survival Analysis with Adversarial Regularization (SAWAR), which is the first method for enabling adversarial robustness to covariate perturbations within a bounded uncertainty set for fully-parametric NN-based SA. We improve calibration and stability on unexpected or noisy data by introducing a loss function based on a Min-Max formulation and Convex Relaxation Based Backward Bounding Pass with Interval Bound Propagation (CROWN-IBP) [20, 21].
The paper is structured as follows: Section II describes the State of the Art; Section III provides background to SA terminology and dataset structure; Section IV provides background on the adversarial and perturbation attacks used; Section V discusses the custom loss employed to deal with those attacks; Section VI provides the results of the experimental activity; finally, Section VII derives conclusions on our findings.
@@ II Related Work
This section focuses on the literature about NNs for SA, with further detail about NN implementations available in a recent survey paper [22]. The first use of NNs in this field was a Feedforward Neural Network (FFNN) to approximate the survival function [23], defined as the probability that no event occurs up to a specific time $t$. More recently, [24] proposed Cox-nnet, a NN that outputs a proportional-hazard [25, 6], which is the assumption that the difference in effect between different instances is constant as a function of time. Building on this idea, [26] developed DeepSurv, a Cox proportional-hazard NN for personalized treatment, which uses dropout layers to improve regularization and reduce overfitting. Instead of proportional hazard estimation, DeepWeiSurv [27] regresses the parameters of a Weibull mixture distribution with NNs.
Several recent works in SA aim to improve model calibration (e.g., by using new terms in the objective function [28, 29, 30]). A calibrated model provides a predicted confidence that is indicative of the true likelihood that the model is correct [31]. Meanwhile, few works have addressed adversarial robustness in SA. In general, a model is adversarially robust if small perturbations to the input do not cause the model to make incorrect predictions [32]. Adversarial robustness and model calibration are loosely coupled: while [33] showed that better calibration does not necessarily improve generalization, [34] showed that incorporating adversarial robustness during training improves trustworthiness in calibration and stability on unexpected data.
Existing adversarial training methods for SA face important tradeoffs between model complexity, performance, and ease of embedding of physical assumptions. Most of these techniques use Conditional Generative Adversarial Networks (CGANs) to generate samples of the time-to-event from a survival distribution in a non-parametric manner [19, 35, 36]. These models (e.g., CGANs) do not assume a functional form and define the survival distribution implicitly [37, 38, 39]. Therefore, there is no mathematical expression that describes the time-to-event Probability Density Function (PDF) value given covariates. One benefit of implicit models is that increased model complexity leads to increased flexibility, which is the ability to fit many other potential data realizations [40]. However, the higher model complexity generally violates the Bayesian Occam’s Razor, which penalizes models for having too much flexibility [40]. To satisfy the Bayesian Occam’s Razor’s “just right“ model complexity, modeling assumptions known as inductive biases are employed in the form of an explicit parametric model [39, 37]. For example, a common inductive bias for SA models is the memoryless property [38], which the exponential distribution satisfies. For this reason, this paper uses the exponential distribution, which is commonly used in SA [41, 42, 43] (and more generally in queuing theory [44, 45], reliability analysis [46, 8], and telecommunications [45]).
@@ III Preliminaries on Survival Analysis
SA [25, 6] is a statistical approach used to analyze and model the time $t\in\mathcal{R_{++}}$ until an event $e\in\{0,1\}$ happens (e.g., a disease occurs, a device breaks). Typically, the study happens over a finite time period and instances/individuals may leave the study early before an event is recorded, with no follow up. In those cases, the dataset records that no event occurred up until the exit time of the study, which is called right censoring [47]. Such an SA dataset consists of $N$ independent instances/individuals, each denoted by index $i$, for which time between entry to a study and the subsequent event is recorded. Upon entry time to a study, the instance’s covariates $\boldsymbol{x}_{i}\in\mathcal{R}^{d}$ (i.e., variables that may affect the event of interest for the study) are measured. In summary, the dataset $\mathcal{D}$ has the following structure:
$\mathcal{D}=\{(\boldsymbol{x_{i}},t_{i},e_{i})\}_{i=1}^{N}.$
@@
III-A Proportional Hazard Model
Let us denote $T$ as a continuous, non-negative, random variable that represents the time-to-event, with PDF $f(t)$ and Cumulative Distribution Function (CDF) $F(t)=P[T\leq t]$. The survival function $S(t):=1-F(t)$ is the complement of this CDF [48, 6].
@@ III-A Proportional Hazard Model
To incorporate the instance’s covariates, we use the proportional hazard model [49, 7], This model assumes that the covariates increase or decrease the hazard by a proportional amount at all times [48, 6],
@@
The exponential proportional hazard model characterizes the baseline hazard rate as a constant function of time, and therefore $\lambda_{0}(t)=\lambda_{0}$. For simplicity, we assume an exponential baseline hazard, and we may “absorb” the $\lambda$ term into the bias of the NN:
@@
The proportional hazard model PDF, CDF, and complement CDF for an exponential proportional hazard model are then given by:
$f_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ $=e^{G_{\theta}(\boldsymbol{x})}e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$
$F_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ $=1.0-e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$
$S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ $=e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}.$
$S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ is also known as the instance survival function.
@@
Since the distribution of the covariates $p(\boldsymbol{x})$ is typically unknown [50], the integral in Eq. 5 is often replaced with a Monte Carlo estimate [28]:
@@ III-B Baseline Objective Functions
Given those definitions, the following loss functions enable finding model parameters $\boldsymbol{\theta}$ from dataset $\mathcal{D}$. In particular, we follow [51, 52, 19] to define two objective functions based on negative log likelihood (Section III-B1) and the rank correlation (Section III-B2).
@@ III-B1 Right Censored Log Likelihood Objective
The $\mathcal{L}_{LL}$ objective is the log of the time-to-event likelihood over $\mathcal{D}$ specified by the proportional hazard model [6]:
@@ III-B2 Ranking Objective
where $\eta(\mathbf{x},\mathbf{y})=\exp(-\frac{(\mathbf{x}-\mathbf{y})}{\sigma})$ and $A_{i,j}$ indicates an “acceptable” comparison pair [51, 52, 53]. We set $\sigma=1$ for all experiments.
@@ III-B3 Combined Objective
The combined right censored negative log likelihood objective and the risk ranking objective, with risk ranking objective weight trade-off $w=\frac{1}{\text{batch\_size}}$ is
@@
This combined objective provides a trade-off between finding the NN parameters $\boldsymbol{\theta}$ that maximize the likelihood of observing $\mathcal{D}$ and maximize a rank correlation between the instance risk scores and the observed failure times. We will denote the combined loss as the baseline objective. Our baseline model’s objective is similar to DRAFT [54] but with a different ranking objective function.
@@
$\boldsymbol{\theta}^{*}=\arg\min_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})$ (10)
In the next section, we address how to develop NN models that are robust against adversarial attacks.
@@ IV Preliminaries on Adversarial Training
Adversarial training [32, 55, 21, 56] is a technique to improve robustness and generalization of NNs to uncertain data. Rather than training on pristine data, adversarial training perturbs the data throughout the training process to encourage the model to optimize for parameters that are resilient against such perturbations. For example, [56] formulated adversarial training as solving a Min-Max robust optimization problem:
@@ IV-A Certifiable Robustness
One such approximation is to use techniques from NN verification during training. For example, CROWN-IBP combines Convex Relaxation Based Backward Bounding Pass (CROWN) and Interval Bound Propogation (IBP) methods to obtain a convex relaxation for the lower bound and upper bound of each layer in the NN with respect to input data [55, 57]. CROWN-IBP is commonly used for verifying NN properties over a range of possible inputs. [58, 55] showed the linear relaxation may be extended to the entire objective function:
@@
This relaxation has several benefits: it is convex with respect to the NN weights, which enables efficient optimization, and by optimizing over an upper bound on the inner maximization, this relaxation represents accounting for a strong adversary during training.
IV-B Adversarial Attacks
Another way to approximate Eq. 19 is to use adversarial attacks. Whereas CROWN-IBP provides an upper bound on the inner maximization (i.e., $\max_{\boldsymbol{\tilde{X}}}\mathcal{L}(\theta;\boldsymbol{\tilde{X}})$), adversarial attack algorithms typically provide a lower bound:
$\max_{\boldsymbol{\tilde{X}}}\ \mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}})\approx\mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}_{FGSM}})$ (13)
Fast Gradient Sign Method (FGSM) [59] uses the gradient of the objective function with respect to the input $\boldsymbol{x}$ to perturb $\boldsymbol{x}$ in the direction that maximizes the objective $\mathcal{L}$. Projected Gradient Descent (PGD) [56] iteratively applies that same type of first-order perturbation $K$ times (or until convergence is reached), but constrains the perturbation to $\boldsymbol{x}$ within a specified set (e.g., $\ell_{p}$-ball). FGSM is a special case of PGD, with $K=1$ and the $\ell_{\infty}$-ball.
@@ IV-C Random Noise
While robustness to a worst-case perturbation is important for many applications, adding random noise to training data has also been shown to improve generalization performance [60]. Therefore, this paper also considers Gaussian Noise perturbations with variance $\epsilon$ bounded within an $\ell_{\infty}$ ball:
@@
We next introduce our approach, SAWAR, to incorporate adversarial regularization with SA models for improved calibration, predictive generalization, and adversarial robustness.
@@ V Survival Analysis with Adversarial Regularization
As discussed in Section I, a key open challenge in SA is in developing models that are robust to uncertainty and variability in the datasets. This paper aims to improve the adversarial robustness of an NN against covariate perturbations within the SA framework. In doing so, this paper proposes the first method for enabling adversarial robustness for fully-parametric NN-based SA, which improves calibration and stability on unexpected or noisy data.
The following sections detail our approach, which extends the baseline objective function in Section III-B3 to include adversarial regularization posed as a Min-Max optimization problem in Section Section IV [21, 55].

### v3

# Survival Analysis with Adversarial Regularization

@@ Abstract
Survival Analysis (SA) models the time until an event occurs, with applications in fields like medicine, defense, finance, and aerospace. Recent work shows that Neural Networks (NNs) can capture complex relationships in SA. However, dataset uncertainties (e.g., noisy measurements, human error) can degrade model performance. To address this, we leverage NN verification advances to create algorithms for robust, fully-parametric survival models. We introduce a robust loss function and use CROWN-IBP regularization to handle computational challenges in the Min-Max problem. Evaluating our approach on SurvSet datasets, we find that our Survival Analysis with Adversarial Regularization (SAWAR) method consistently outperforms baselines under various perturbations with respect to Negative Log Likelihood (NegLL), Integrated Brier Score (IBS), and Concordance Index (CI). This demonstrates that adversarial regularization enhances SA performance and calibration, mitigating data uncertainty and improving generalization across diverse datasets up to 150% across all perturbation magnitudes.
@@
Survival Analysis, Adversarial Robustness, Neural Networks, Calibration, CROWN-IBP, Cox-Weibull
@@ I Introduction
Survival Analysis (SA) models the time until an event occurs and has extensive applications. For example, SA has been used extensively in medicine for evaluating treatments in cancer patients and drug trials [1]. Early methods relied on linear parametric models [2], but recent advancements utilize Neural Networks (NNs) to handle complex time-to-event data [3].
A major challenge in Survival Analysis (SA) is developing models robust to dataset uncertainty and variability. This is particularly essential for clinical applications where treatment efficacy must be demonstrated across diverse patient demographics, institutions, and data acquisition settings [4, 5, 6]. Data uncertainty can arise from human errors, such as poor handwriting by physicians [7] and ineffective medical record-keeping [8], as well as inherent noise during data collection [9]. Many approaches assume perfect measurement of covariates, neglecting aleatoric uncertainty [10], which in SA is associated with noise in the log of measured time-to-event [11]. Further, adversaries can manipulate data to make models over- or under-confident [12]. Despite these issues and the sensitivity of NNs to input uncertainties, robustness techniques against data perturbations in SA remains underexplored.
@@
  • Survival Analysis with Adversarial Regularization (SAWAR) for training robust, fully-parametric, NN-based SA models using adversarial covariate perturbations and convex relaxations, enhancing model calibration and stability against noisy data.
  • Experiments demonstrating the sensitivity of existing NN-based SA methods to covariate perturbations across 10 publicly available medical SA datasets, highlighting an area for improvement in the current state of the art.
@@ II Related Work
This section reviews NN literature for SA, with more details in [13]. The first use of NNs in SA was a shallow Feedforward Neural Network (FFNN) to approximate the survival function [14]. Later, deeper models were developed, such as Cox-nnet [15], a NN for Cox Proportional Hazard (CPH) modeling [11]. Building on this, DeepSurv [16] introduced a CPH NN with dropout layers for better generalization in personalized treatment. Then, mixture models such as DeepWeiSurv [17] further advanced the field by using NNs to regress Weibull mixture distribution parameters. These methods enhance modeling capability but do not always improve model calibration and robustness.
Recent works focus on improving model calibration [18, 19, 20], which ensures predicted confidence matches true likelihood [21]. However, few works address adversarial robustness in SA, where small input perturbations should not mislead the model [22]. Although better calibration does not always enhance generalization [23], integrating adversarial robustness improves model trustworthiness in calibration and and stability on unexpected data [24].
Existing adversarial training methods for SA involve tradeoffs between model complexity, performance, and incorporating physical assumptions. Most techniques use Conditional Generative Adversarial Networks (CGANs) to generate time-to-event samples from a survival distribution non-parametrically [25, 26, 27]. These models implicitly define the survival distribution without assuming a functional/analytical form, offering increased flexibility to fit various data realizations [28, 29, 30]. However, this flexibility can violate Bayesian Occam’s Razor, which penalizes overly complex models [30]. To balance complexity and performance, inductive biases are used in explicit parametric models [29]. One common bias in SA is the memoryless property, satisfied by the exponential distribution [28]. Thus, this paper employs the exponential CPH model, widely used in SA [31, 32, 33], as well as in queuing theory [34], reliability analysis [3], and telecommunications [35].
@@ III Preliminaries on Survival Analysis
SA [11] is a statistical approach used to analyze and model the time $t\in\mathcal{R_{++}}$ until an event $e\in\{0,1\}$ happens (e.g., a disease occurs, a device breaks). Typically, the study happens over a finite time period and instances/individuals may leave the study early before an event is recorded, with no follow up. In those cases, the dataset records that no event occurred up until the exit time of the study, which is called right censoring [36]. Such an SA dataset consists of $N$ independent instances/individuals, each denoted by index $i$, for which time between entry to a study and the subsequent event is recorded. Upon entry time to a study, each instance’s covariates $\boldsymbol{x}_{i}\in\mathcal{R}^{d}$ (i.e., variables that may affect the event of interest for the study) are measured. In summary, the dataset $\mathcal{D}$ has the following structure:
$\mathcal{D}=\{(\boldsymbol{x}_{i},t_{i},e_{i})\}_{i=1}^{N}.$ (1)
@@
III-A Cox Proportional Hazard Model
Let us denote $T$ as a continuous, non-negative, random variable that represents the time-to-event, with Probability Density Function (PDF) $f(t)$ and Cumulative Distribution Function (CDF) $F(t)=P[T\leq t]$. The survival function $S(t):=1-F(t)$ is the complement of this CDF [37, 11].
@@ III-A Cox Proportional Hazard Model
To incorporate the instance’s covariates, we use the prevalent CPH model [2]:
@@
The exponential CPH model characterizes the baseline hazard rate as $\lambda_{0}(t)=\lambda_{0}$. For simplicity, we assume an exponential baseline hazard and “absorb” the $\lambda$ term into the bias of the NN:
@@
The exponential CPH model’s PDF, CDF, and complement CDF are $f_{\boldsymbol{\theta}}(t|\boldsymbol{x})=e^{G_{\theta}(\boldsymbol{x})}e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$, $F_{\boldsymbol{\theta}}(t|\boldsymbol{x})=1.0-e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$ , $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})=e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$, respectively. $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ is also known as the instance survival function.
@@
Since the distribution of the covariates $p(\boldsymbol{x})$ is typically unknown, the integral in Eq. 6 is often replaced with a Monte Carlo estimate [18]:
@@
Given these definitions, the next section describes the baseline objective for finding model parameters $\boldsymbol{\theta}$ from dataset $\mathcal{D}$.
@@ III-B Baseline Objective Functions
We follow [38, 25] to define two objective functions based on negative log likelihood (Section III-B1) and the rank correlation (Section III-B2).
@@ III-B1 Right Censored Log Likelihood Objective
The $\mathcal{L}_{LL}$ objective is the log of the time-to-event likelihood over $\mathcal{D}$ specified by the CPH model [11]:
@@ III-B2 Ranking Objective
where $\eta(\mathbf{x},\mathbf{y})=\exp(-\frac{(\mathbf{x}-\mathbf{y})}{\sigma})$ and $A_{i,j}$ indicates an “acceptable” comparison pair [38, 39]. We set $\sigma=1$ for all experiments.
@@ III-B3 Combined Objective
The combined right censored negative log likelihood objective and the risk ranking objective, with risk ranking objective weight trade-off $w=\frac{1}{\text{batch\_size}}$, is
@@
This combined objective provides a trade-off between finding the NN parameters $\boldsymbol{\theta}$ that maximize the likelihood of observing $\mathcal{D}$ and maximize a rank correlation between the instance risk scores and the observed failure times. We will denote the combined loss as the baseline objective. Our baseline model’s objective is similar to Deep Regularized Accelerated Failure Time (DRAFT) [25] but with a different ranking objective function. Thus, we consider the baseline training method for robust SA baseline method as DRAFT.
@@
$\boldsymbol{\theta}^{*}=\arg\min_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e}).$ (11)
In the following sections, we address how to develop NN models that are robust against adversarial perturbations.
@@ IV Preliminaries on Adversarial Training
Adversarial training [22, 40, 41] improves robustness and generalization of NNs to uncertain data. Rather than training on pristine data, most adversarial training approaches perturb the data throughout the training process to encourage the model to optimize for parameters that are resilient against such perturbations. For example, [42] formulated adversarial training as solving a Min-Max robust optimization problem:
@@ IV-A Certifiable Robustness
One such approximation is to use techniques from NN verification during training. For example, CROWN-Interval Bound Propagation (CROWN-IBP) combines CROWN and Interval Bound Propogation (IBP) methods to obtain a convex relaxation for the lower bound and upper bound of each layer in the NN with respect to input data [40, 43]. CROWN-IBP is commonly used for verifying NN properties over a range of possible inputs. [44, 40] showed the linear relaxation can be extended to the entire objective function:
@@
This relaxation is convex with respect to NN weights, enabling efficient optimization, and by optimizing over an upper bound on the inner maximization, it accounts for a strong adversary during training.
IV-B Adversarial Perturbations
Another way to approximate Eq. 20 is to use adversarial perturbations. Whereas CROWN-IBP provides an upper bound on the inner maximization (i.e., $\max_{\boldsymbol{\tilde{X}}}\mathcal{L}(\theta;\boldsymbol{\tilde{X}})$), adversarial perturbation algorithms typically provide a lower bound:
$\max_{\boldsymbol{\tilde{X}}}\ \mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}})\geq\mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}}_{FGSM}).$ (14)
Fast Gradient Sign Method (FGSM) [45] uses the gradient of the objective function with respect to the input $\boldsymbol{x}$ to perturb $\boldsymbol{x}$ in the direction that maximizes the objective $\mathcal{L}$. Projected Gradient Descent (PGD) [42] iteratively applies that same type of first-order perturbation $K$ times (or until convergence is reached), but constrains the perturbation of $\boldsymbol{x}$ to remain within a specified set (e.g., $\ell_{p}$-ball). FGSM is a special case of PGD, with $K=1$ and the $\ell_{\infty}$-ball.
@@ IV-C Random Noise
While robustness to a worst-case perturbation is important for many applications, adding random noise to training data has also been shown to improve generalization performance [46]. Therefore, this paper also considers Gaussian Noise perturbations with variance $\epsilon$ bounded within an $\ell_{\infty}$-ball:
@@
Building on this foundation, we next introduce our approach, SAWAR, that develops robust SA models based on adversarial regularization.
@@ V Survival Analysis with Adversarial Regularization
This paper addresses the open challenge in SA of developing models robust to dataset uncertainty and variability, as motivated in Section I. We propose a method for adversarial robustness in fully-parametric NN-based SA, enhancing calibration and stability on unexpected or noisy data. Our approach extends the objective function (Eq. 10) with adversarial regularization as a Min-Max optimization problem (Eq. 12) [41, 40], as detailed in the following sections.